A father is now 28 years older than his son. Ten years ago the father was 15 times as old as his son. How old is the father now?
Answer by fractalier(6550)

Call the father and son's age NOW, f and s. Then we have
f = s + 28 and
f - 10 = 15(s - 10)
Substitute the first into the second equation and get
s + 28 - 10 = 15(s - 10)
s + 18 = 15s - 150
168 = 14s
s = 12
f is then 40